# Separation and collapse of equilibria inequalities on AND–OR trees without shape constraints

### v2

@@ 2 Definitions and Preliminaries
We observe a randomized directional algorithm in the broad sense that is not an RDA. Let $T$ be a complete binary tree of height 2 such that the root is labeled with AND, and each of child nodes $v_{1},v_{2}$ of the root is labeled with OR. Let $x_{1},x_{2}$ be the leaves that are child nodes of $v_{1}$, and $x_{3},x_{4}$ be the leaves that are child nodes of $v_{2}$. Let $\alpha\in\mathcal{A}_{\mathrm{dir}}(T)$ be the algorithm whose priority of probing leaves is in the order of $x_{1},x_{2},x_{3},$ and $x_{4}$. Let $\beta\in\mathcal{A}_{\mathrm{dir}}(T)$ be the algorithm whose priority of probing leaves is in the order of $x_{2},x_{1},x_{4},$ and $x_{3}$. Let $A\in\mathcal{D}(\mathcal{A}_{\mathrm{dir}}(T))$ be the randomized algorithm such that $A$ works as $\alpha$ with probability $1/2$ and as $\beta$ with probability $1/2$. Let “$x_{i}\to x_{j}$” denote the event that $x_{i}$ is probed before $x_{j}$. Then we have the following, where $\mathrm{Pr}_{A}$ denotes probability with respect to $A$.
@@
Suppose that $T$ is as above.
@@
    Saks-Wigderson’s $d$ is defined as follows [14].
@@
    $R_{\mathrm{dir},i}(T):=$ $\min_{A\in\mathcal{D}(\mathcal{A}_{\mathrm{DF}}(T))}~{}\max_{\omega\in\Omega_{i}(T)}\mathrm{cost}(A,\omega)$ (2.6)
@@
It is immediate that $R(T)\leq R_{\mathrm{DF}}(T)\leq R_{\mathrm{dir}}(T)\leq d(T)$. Under a certain hypothesis on $T$ (weakly balanced in the sense of Definition 1.1), it is shown that $R(T)=d(T)$ (Theorem 1.2). In this paper, we are going to show $R_{\mathrm{DF}}(T)=d(T)$ regardless of the assumption of weakly balanced.
@@
(2). The inequality $\geq$ is obvious. Let $S_{*}$ be an element of $\mathcal{D}(\Omega^{\prime})$ such that $\mathrm{cost}(A,S_{*})=\max_{S\in\mathcal{D}(\Omega^{\prime})}\mathrm{cost}(A,S)$. Let $\omega_{*}$ be an element of $\Omega^{\prime}$ such that it maximizes $\mathrm{cost}(A,\omega)$ among all $\omega\in\Omega^{\prime}$ such that $S(\omega)>0$. Then $\mathrm{cost}(A,S_{*})=\sum_{\omega\in\Omega^{\prime}}S(\omega)\mathrm{cost}(A,\omega)\leq\mathrm{cost}(A,\omega_{*})\leq\max_{\omega\in\Omega^{\prime}}\mathrm{cost}(A,\omega)$. Thus the inequality $\leq$ holds.
@@ 3.1 Replacement of sub-algorithms
The following result is written in [14] without a proof, and we include it here.
@@
[14, Lemma 3.1] There exists an RDA on $T$ that attains both $d_{0}(T)$ and $d_{1}(T)$.
@@ 5 Concluding remarks
In [5], it is shown that if an AND–OR tree $T$ satisfies $R(T)=d(T)$, there are infinitely many randomized algorithms which achieve $R(T)$. By our result, the assumption $R=d$ is equivalent to $R=R_{\mathrm{DF}}$.

### v3

@@ 2 Definitions and Preliminaries
(A depth-first algorithm that is not directional) We observe a depth-first algorithm that is not directional. Let $T$ be a complete binary tree of height 2 such that the root is labeled with AND, and each of child nodes $v_{1},v_{2}$ of the root is labeled with OR. Let $x_{1},x_{2}$ be the leaves that are child nodes of $v_{1}$, and $x_{3},x_{4}$ be the leaves that are child nodes of $v_{2}$. Let $\alpha\in\mathcal{A}_{\mathrm{DF}}(T)$ be the algorithm whose priority of probing leaves is as follows.
  • $\alpha$ begins with probing $x_{1}$.
  • If $x_{1}$ has value $1$ then $\alpha$ probes $x_{3}$. If $x_{3}$ has value $1$ then $\alpha$ returns $1$ and halts. Otherwise, $\alpha$ returns value of $x_{4}$ and halts.
  • If $x_{1}$ has value 00 then $\alpha$ probes $x_{2}$. If $x_{2}$ has value 00 then $\alpha$ returns 00 and halts. Otherwise, $\alpha$ probes $x_{4}$. If $x_{4}$ has value $1$ then $\alpha$ returns $1$ and halts. Otherwise, $\alpha$ returns value of $x_{3}$ and halts.
For truth assignment $\omega_{1}=(1,1,0,0)$ (that is, $\omega_{1}$ assigns $(1,1,0,0)$ to $(x_{1},x_{2},x_{3},x_{4})$), $\alpha$ probes $T$ in the order of $x_{1},x_{3},$ and $x_{4}$. For truth assignment $\omega_{2}=(0,1,0,0)$, $\alpha$ probes $T$ in the order of $x_{1},x_{2},x_{4}$ and $x_{3}$. Since priority of $x_{3}$ and $x_{4}$ is not fixed, $\alpha$ is not directional.
It holds that $\alpha\in\mathcal{A}_{\mathrm{DF}}(T)\setminus\mathcal{A}_{\mathrm{dir}}(T)$. At the same time, we have $)\alpha\in\mathcal{D}(\mathcal{A}_{\mathrm{DF}}(T))\setminus\mathcal{D}(\mathcal{A}_{\mathrm{dir}}(T)$.
Example 2.2.
(A randomized directional algorithm in the broad sense that is not an RDA) We observe a randomized directional algorithm in the broad sense that is not an RDA. Let $T$ be as in Example 2.1. Let $\alpha\in\mathcal{A}_{\mathrm{dir}}(T)$ be the algorithm whose priority of probing leaves is in the order of $x_{1},x_{2},x_{3},$ and $x_{4}$. Let $\beta\in\mathcal{A}_{\mathrm{dir}}(T)$ be the algorithm whose priority of probing leaves is in the order of $x_{2},x_{1},x_{4},$ and $x_{3}$. Let $A\in\mathcal{D}(\mathcal{A}_{\mathrm{dir}}(T))$ be the randomized algorithm such that $A$ works as $\alpha$ with probability $1/2$ and as $\beta$ with probability $1/2$. Let “$x_{i}\to x_{j}$” denote the event that $x_{i}$ is probed before $x_{j}$. Then we have the following, where $\mathrm{Pr}_{A}$ denotes probability with respect to $A$.
@@
Suppose that $T$ is an AND-OR tree in the sense of Convention 1.1.
@@
    Saks-Wigderson’s $d$ is defined as follows [SW86].
@@
    $R_{\mathrm{dir},i}(T):=$ $\min_{A\in\mathcal{D}(\mathcal{A}_{\mathrm{dir}}(T))}~{}\max_{\omega\in\Omega_{i}(T)}\mathrm{cost}(A,\omega)$ (2.5)
@@
(2). The inequality $\geq$ is obvious. Let $S_{*}$ be an element of $\mathcal{D}(\Omega^{\prime})$ such that
$\mathrm{cost}(A,S_{*})=\max_{S\in\mathcal{D}(\Omega^{\prime})}\mathrm{cost}(A,S).$
Let $\omega_{*}$ be an element of $\Omega^{\prime}$ such that it maximizes $\mathrm{cost}(A,\omega)$ among all $\omega\in\Omega^{\prime}$ such that $S_{*}(\omega)>0$. Then $\mathrm{cost}(A,S_{*})=\sum_{\omega\in\Omega^{\prime}}S_{*}(\omega)\mathrm{cost}(A,\omega)\leq\mathrm{cost}(A,\omega_{*})\leq\max_{\omega\in\Omega^{\prime}}\mathrm{cost}(A,\omega)$. Thus the inequality $\leq$ holds.
@@ 3.1 Replacement of sub-algorithms
The following result is written in [SW86] without a proof, and we include it here.
@@
[SW86, Lemma 3.1] There exists an RDA on $T$ that attains both $d_{0}(T)$ and $d_{1}(T)$.
@@ 5 Concluding remarks
In [I25], it is shown that if an AND–OR tree $T$ satisfies $R(T)=d(T)$, there are infinitely many randomized algorithms which achieve $R(T)$. By our result, the assumption $R=d$ is equivalent to $R=R_{\mathrm{DF}}$.